Week of 161003

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Maarten + Luca (SCODs), Maria + Julia (WLCG), Ivan G. (ATLAS), Alberto A.(Monitoring), Gavin (Computing), Jesus (Storage), Vincent (Security), Andrea M. (MW), David C.(DB), Marian (Network)
  • remote: Antonio (LHCb + CNAF), Stefano (CMS), FaHui (ASGC), Eric (BNL), Vincenzo (EGI), David (FNAL), Rolf (IN2P3), Victor (JINR), Sang-Un (KISTI), Gareth (RAL), Christian (NDGF), Dennis (NL-T1), Kyle (OSG), Pepe (PIC), Di (TRIUMF)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities and global report
      • Grid is running full capacity, production continues with the new Sherpa MC samples.
      • Productive Software and Computing week last week.
    • Problems:
      • EOSATLAS service was degraded September 27th night (high latencies between diskservers and the EOSATLAS headnode), stabilized one day later, shortage of disk space at Meyrin, geotagging disabled (IT ticket OTG:0033142).

  • CMS reports ( raw view) -
    • Running at capacity
    • Widespread xrootd failures during the weekend traced to one user submitting a particular analysis workflow that was causing zillions of file open attempts at the same time crashing redirectors. Jobs eventually killed by CMS analysis support, user contacted, work in progress by experts and admins to put a limit to # of simultaneous file open attempts by each client.
      • WAN data access started as a fallback for local data read failure, but now we can't live w/o it.
    • Otherwise quiet week, business as usual

  • ALICE -
    • KISTI incident:
      • Users reported persistent job failures due to CVMFS errors on Sep 26
      • The CVMFS Squid servers had /var full due to excessive logging
      • A cleanup was performed, but the services were accidentally left down
      • Yet more failures were reported on Sep 30
      • The Squid services were then restarted that evening, thanks!
      • The job success rates improved dramatically as a result

Sites / Services round table:

  • ASGC: NTR
  • BNL: NTR
  • CNAF: The mandatory security update of last Friday on canl-c broke some configuration on our CEs causing problems for CMS and LHCb. the site is investigating the issue.
  • EGI: refer to the security entry
  • FNAL: NTR
  • GridPP:
  • IN2P3: NTR
  • JINR: No issues for the week, CMS requested a AAA update on Friday evening.
  • KISTI: downtime on Friday for router intervention, issues during the weekend to contact CERN and job profile went to 0.
  • KIT:
  • NDGF: After the switch swap Thursday, we have had issues with routing from the Norwegian pools over IPv4. Link dies every 5 min. Uninett and Nordunet are investigating.
  • NL-T1:
  • NRC-KI:
  • OSG: NTR
  • PIC: NTR
  • RAL: We are planning to replace the network router that manages our external data flows on Wednesday (5th Oct). This will interrupt external data access but not other services.
  • TRIUMF:NTR

  • CERN computing services: NTR
  • CERN storage services:
    CASTOR UPDATES
    ATLAS + PUBLIC: Tuesday 4th October from 10AM to noon.
    LHCB + CMS + ALICE: Wednesday 5th October from 10AM to noon.
  • CERN databases: NTR
  • GGUS: release deployed on 28th September

  • Monitoring:
    • Link to the monitoring portal: http://monit.cern.ch
    • Draft reports for the 09 2016 availability were sent around today
    • LHCb VO feed doesn’t contain storage elements GGUS:124027

  • MW Officer: There is an issue with ARC-CE + HTCondor >= 8.5.5 while querying for running jobs. HTCondor changed the default condor_q output such that only the current user's jobs are returned. GGUS:123947 and http://bugzilla.nordugrid.org/show_bug.cgi?id=3603 report a workaround to fix this. The problem has been reported last week by LHCb, but it affects also other VOs.
  • Networks: NTR
  • Security: canl-c security advisory, sites should update their worker nodes as well.

AOB:

Edit | Attach | Watch | Print version | History: r14 < r13 < r12 < r11 < r10 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r14 - 2016-10-03 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback